OpenStudy (anonymous):

What is the range of y = 5x -2 if the domain is {-3, -1, 0, 1, 3}? {-13, -3, 2, 3, 13} {12, 3, 0, 3, 13} {-17, -7, -2, 3, 13}

OpenStudy (anonymous):

@BunnyBree

OpenStudy (anonymous):

HEEELLLPPP

OpenStudy (texaschic101):

domain is all the x values and range is all the y values we are given the domain (x values) so sub in the x values to find the y values (the range) y = 5x - 2.....when domain (x value) is -3 y = 5(-3) - 2 y = -15 - 2 y = -17...so when the domain is -3, the range is -17 next number y = 5x - 2....when domain (x) is -1 y = 5(-1) - 2 y = -5 - 2 y = -7 .... so when the domain is -1, the range is -7 you see how I am doing this ?

OpenStudy (anonymous):

kind of

OpenStudy (anonymous):

so the answer would be C

OpenStudy (texaschic101):

just sub in your domain (x values) into the equation to find the rage (y values) answer is C

OpenStudy (anonymous):

ok thanks I have one more question

OpenStudy (anonymous):

What is the solution set for the inequality 3x - 7 >9 if the replacement set is {-2, 0, 5, 7, 11, 13}? {-2, 0, 5} {7, 11, 13} {-2, 0, 5, 7, 11, 13} {0, 5, 7, 11}